OpenStudy (anonymous):

A new cruise ship line has just launched 3 new ships: The pacific paradise, the caribbean paradise, and the mediterranean paradise. The caribeean paradise has 38 more deluxe staterooms than the pacidic paradise. The mediterranean Paradise has 40 fewer deluxe staterooms than four times the number of deluxe staterooms on the pacific paradise. Find the number of deluxe staterooms for each of the ships if the total number of deluxe staterooms for the three ships is 814.

OpenStudy (anonymous):

let p, c and m be the the number of deluxe staterooms on the pacific paradise, caribbean paradise and the mediterranean paradise, respectively. Solve the following equations for p, c and m:\[\{c\text{ = }p+38,m=-40+4p,c+p+m=814\}\]\[\{c=174,p=136,m=504\} \]